For the equation x-(sqrt)x=0, solve for all values of x that satisfies the equation and show the work.
Answer by ptaylor(2198)

To solve:
(1) x-sqrt(x)=0
We want to get rid of the radical. We'll do this by, first, isolating it on the left side of the equation and then we'll square both sides.
subtracting x from both side of the equation, we get:
x-x-sqrt(x)=-x simplifying we have:
-sqrt(x)=-x now we'll square both sides:
(-sqrt(x))^2=(-x)^2 and we get:
x=x2 subtract x from both sides:
x2-x=0 or
x(x-1)=0 and
x=0
x=1
ck by substituting into (1)
0-0=0
1-1=0
Hope this helps.---Happy holidays.-----ptaylor